# Algebraic Control: Complete Stable Inversion with Necessary and Sufficient Conditions

### v3

@@ Abstract
In this paper, we establish necessary and sufficient conditions for stable inversion, addressing challenges in non-minimum phase, non-square, and singular systems. An $\mathcal{H}_{\infty}$-based algebraic approximation is introduced for near-perfect tracking without preview. Additionally, we propose a novel robust control strategy combining the nominal model with dual feedforward control to form a feedback structure. Numerical comparison demonstrates the approach’s effectiveness.
@@ I Introduction
The study of stable inversion began with Brockett’s introduction of functional reproducibility in 1965 [2]. Silverman extended these concepts to multivariable systems in 1969 [3], followed by geometric formulations introduced by Basile and Marro [4] and applications in reduced-order control by Moylan [5]. The stable inversion in discrete-time systems was first addressed by Tomizuka [6], while Hunt explored non-causal inversion techniques [7].
@@ II-A Algebraic Preliminaries
where $U_{R}(s)\in\mathbb{R}[s]^{n_{y}\times n_{y}}$ and $V_{R}(s)\in\mathbb{R}[s]^{n_{u}\times n_{u}}$ are $\mathcal{R}$-Unimodular matrices, $\Lambda(s)=\text{diag}[d_{1}(s)\ \dots\ d_{r}(s)]$ with unique monic $d_{i}(s)\in\mathcal{R}$ such that $d_{i}$ divides $d_{i+1}$. The matrix $S_{m}(s)\in\mathbb{R}[s]^{n_{y}\times n_{u}}$ is called as Smith Normal form of $\grave{P}$.
Let $V$ be a vector space over the field $\mathbb{R}(s)$ with dimension $k$, consisting of $n$-tuples such that a basis of vector polynomials can always be found for $V$ [page 22, [22]]. A minimal basis of $V$ is defined as a $k\times n$ polynomial matrix $P_{m}$ [20]. Adapted Forney’s Theorem [20, Section 3(4.)] states that if $y=xP_{m}$ is a polynomial $n$-tuple, then $x$ must be a polynomial $k$-tuple.
For the rank notation, akin to [23], consider $P$ belongs to $\mathbb{R}(s)^{n_{y}\times n_{u}}$. Then, the rank of $P$ is defined as the maximum size of any linearly independent subset of its columns in the field $\mathbb{R}(s)$, denoted by $\text{rank}_{\mathbb{R}(s)}(P)$, where $\text{rank}_{\mathbb{R}(s)}(P)\neq\text{rank}_{\mathbb{R}}(P)$.
Suppose $\text{rank}_{\mathbb{R}(s)}(P)=r$, where $1\leq r\leq\min(n_{u},n_{y})$. Define $J=\{j_{1},\dots,j_{r}\}\subseteq\{1,\dots,n_{\color[rgb]{0,0,0}u}\}$ as an ordered indexed set corresponding to $P(s)$’s linearly independent columns. We then define the matrix $L$ as
@@
Some further notations throughout the paper are: (.$\|(.)(t)\|_{\infty}\triangleq\text{ess sup}_{t}|(.)(t)|$, any complex number can be expressed as (.)±$\Re(.)\pm j\Im(.)\in\mathbb{C}$, $\sigma$ represents singular values. Time domain square-integrable functions are denoted by $\mathcal{L}_{2}(-\infty,\infty)$. Its causal subset is given by $\mathcal{L}_{2}[0,\infty)$. For the frequency domain, including at $\infty$, $\mathcal{L}_{2}(j\mathbb{R})$ represents square-integrable functions on $j\mathbb{R}$, and $\mathcal{L}_{\infty}(j\mathbb{R})$ denotes bounded functions on $\Re(s)=0$. All these functions in Lebesgue spaces may be either matrix-valued or scalar. Then, $\mathcal{R}\mathcal{H}_{\infty}$ denotes the set of real rational $\mathcal{L}_{\infty}(j\mathbb{R})$ functions analytic in $\Re(s)>0$, $(.)^{\dagger}$ represents pseudo-inverse yielding unit matrix under multiplication, $(.)^{*}$ denotes complex conjugates transpose, and .,$\langle$.,$\rangle$ . , . ⟩ denotes the inner product. Now, among various definitions for multivariable zeros of a Transfer Function Matrix (TFM), $P(s)$, we adopt the definition given in [24, chapter 4.5.3] as
@@ II-B Problem Statement
where $A\in\mathbb{R}^{n\times n}$, $B\in\mathbb{R}^{n\times n_{u}}$, $C\in\mathbb{R}^{n_{y}\times n}$, $x,x_{0}\in\mathbb{X}\subseteq\mathbb{R}^{n},x(t)\in\mathcal{L}_{2}[0,\infty),y\in\mathbb{Y}\subset\mathbb{R}^{n_{y}},y(t)\in\mathcal{L}_{2}[0,\infty),u\in\mathbb{U}\subset\mathbb{R}^{n_{u}}$, and $u(t)\in\mathcal{L}_{2}[0,\infty)$. Moreover, the TFM representation of $\Xi(u,0)\triangleq C(sI-A)^{-1}B=P(s)$ and since all real rational strictly proper transfer matrices with no poles on the imaginary axis form $\mathcal{RL}_{2}(j\mathbb{R})\subset\mathcal{L}_{2}(j\mathbb{R})$ [25, page 48], $P(s)$ belongs to $\mathcal{L}_{2}(j\mathbb{R})\cap\mathbb{R}(s)^{n_{y}\times n_{u}}$. The system, $\Xi$, is; P1) a non-minimum phase s.t. $Z_{R}\neq\emptyset$, P2) either $n_{y}=n_{u}$ (square) or $n_{y}\neq n_{u}$ (nonsquare), and causal ($t>0$).
@@
Under Assumption 1, let $\Xi$ denote the forward system given in (6), and let $y_{1}(t)$ be an observed output trajectory. We define the right inverse $\Xi_{i}^{\Gamma}:\mathbb{Y}\times\mathbb{X}\to\mathbb{U}$, which, given $y_{1}(t)$ and a prescribed inverse system’s initial states, $\bar{x}_{0}\in\mathbb{X}$, produces the control input
$u_{inv}(t)=\Xi_{i}^{\Gamma}(y_{1}(t),\bar{x}_{0}),\quad\|u_{inv}(t)\|_{\infty}<\infty{\color[rgb]{0,0,0}.}$ (7)
@@
The terminology here (cf. [26]) also lets $y_{1}(t)$ denote a desired trajectory [10]. With Assumption 1 and $\|e\|_{\infty}<\infty$, (8) is Lyapunov-stable. If exact inversion ($\alpha=\varphi=0$) is unattainable, we can employ a robust stabilising approximate inverse—one viable choice among the many approximations acknowledged in [11]—with an irreducible error $\varphi$.
@@
As is typical in exact stable inversion, we initially assume that $(A,B,C)$ in (6) and the initial states $x_{0}$ and $\bar{x}_{0}$ are known. In Section III-D, this assumption is relaxed to accommodate model uncertainty, and in Sections III-C-III-D, the requirement on $x_{0}$ and $\bar{x}_{0}$ is fully removed, thus allowing unknown initial states but resulting in only an approximate solution. For further details on initial states, see [12].
To solve Problem 1.a algebraically, define $y_{ic}(t)\triangleq\Xi(0,x_{0})$, $y\triangleq y_{1}-y_{ic}$, $u_{inv}^{(0)}(t)\triangleq\Xi_{i}^{\Gamma}(0,\bar{x}_{0})$, and $u(t)\triangleq\Xi_{i}^{\Gamma}(y(t),0)$ to represent the trajectories and control inputs for known or zero initial states. Then, Remark 2 let us re-define (7) as $u_{inv}(t)\triangleq\Xi_{i}^{\Gamma}(y_{1}(t),\bar{x}_{0})-u_{inv}^{(0)}(t)$. Applying the Laplace transform (see Paley–Wiener Th. [27, p. 104] for existence) yields
$\mathscr{L}\{y_{1}(t)-{y}_{ic}(t)\}={C}(s{I}-{A})^{-1}{B}U(s)=Y(s)$ (10)
@@
From this point on, similar to the [19, Theorem 8.5.2], Problem 1.a reduces to algebraically solving the rational matrix equation where $\|u(t)\|_{\infty}<\infty\iff\|u_{inv}(t)\|_{\infty}<\infty$.
@@ III-A Right Inverse
where $U(s)\in\mathbb{R}(s)^{n_{u}}$, and $\grave{Y}(s)\in\mathbb{R}{\color[rgb]{0,0,0}(}s{\color[rgb]{0,0,0})}^{n_{y}}$. Then, using the invariant factor theorem given by (II-A) on $\grave{P}(s)$ in (13), we have
$U(s)$ $=[\hat{V}_{R1}\ \hat{V}_{R2}]\begin{bmatrix}\Lambda^{-1}(s)&\textbf{0}_{(n_{u}-r)\times r}\\ \textbf{0}_{(n_{u}-r)\times r}&\textbf{0}_{(n_{u}-r)\times(n_{y}-r)}\end{bmatrix}\begin{bmatrix}\hat{U}_{R1}\\ \hat{U}_{R2}\end{bmatrix}\grave{Y}(s)$
@@
$=\Biggr{(}[\hat{V}_{R1}\ \hat{V}_{R2}]\begin{bmatrix}\Lambda^{-1}(s)&\textbf{0}_{(n_{u}-r)\times r}\\ \textbf{0}_{(n_{u}-r)\times r}&\textbf{0}_{(n_{u}-r)\times(n_{y}-r)}\end{bmatrix}\begin{bmatrix}\hat{U}_{R1}\\ \hat{U}_{R2}\end{bmatrix}$
@@ III-B Algebraic Necessary and Sufficient Conditions
(2:) Note that $(I_{n_{u}}-\hat{V}_{R1}{V}_{R1})$ in (III-A) is pure polynomial and $\kappa\in\mathcal{R}\mathcal{H}_{\infty}$. Therefore, $(I_{n_{u}}-\hat{V}_{R1}{V}_{R1})\kappa$ does not contribute any unstable solution(s) to $U(s)$. So, considering only the $\hat{V}_{R1}(s)\Lambda^{-1}(s)\hat{U}_{R1}(s)\grave{Y}(s)$ is enough for stability. Lets consider two cases on $\grave{P}(s)$; C1: $\operatorname*{rank}_{\mathbb{R}(s)}(\grave{P})=\min(n_{u},n_{y})$, C2: $\operatorname*{rank}_{\mathbb{R}(s)}(\grave{P})<\min(n_{u},n_{y})$. For C2, re-writing (III-A) and ignoring $(I_{n_{u}}-\hat{V}_{R1}{V}_{R1})\kappa$, we get
@@
where there always exist $\hat{V}^{\ell}_{R1}$ satisfying $\hat{V}^{\ell}_{R1}\hat{V}_{R1}=I_{r}$. Now assume that either $\grave{P}(s)$ or $U_{R1}\Lambda\hat{V}_{R1}^{\color[rgb]{0,0,0}\ell}$ is not a minimal basis. The minimal basis $\grave{P}_{m}(s)$ of an $n_{y}$-dimensional vector space of $n_{u}$-tuples over $\mathbb{R}[s]$ for (13) is given by $\grave{P}_{m}(s)=L_{T}(s)\grave{P}(s)$ for condition C1, and $\grave{P}_{m}(s)=L_{T}(s)U_{R1}\Lambda\hat{V}_{R1}^{\color[rgb]{0,0,0}\ell}$ for condition C2. Here, the $\mathcal{R}$-Unimodular transformation $L_{T}(s)\in\mathbb{R}{\color[rgb]{0,0,0}[}s{\color[rgb]{0,0,0}]}^{n_{y}\times n_{y}}$, which yields the minimal basis, exists with constant (degree zero) determinant (see [20, Remark 2]). Then, left multiplying $L_{T}(s)$ to (13) or (16) results in
@@
Now define ${\color[rgb]{0,0,0}\grave{Y}}(s)\triangleq\frac{1}{\Delta_{\color[rgb]{0,0,0}Y}}{{\color[rgb]{0,0,0}\grave{Y}_{N}}}(s)$ such that ${\color[rgb]{0,0,0}\grave{Y}_{N}}(s)\in\mathbb{R}[s]^{n_{y}}$, and the scalar $\Delta_{\color[rgb]{0,0,0}Y}\in\mathbb{R}[s]$. So, multiplying (17) by $\Delta_{\color[rgb]{0,0,0}Y}$ yields
$\grave{P}_{m}(s)\Delta_{\color[rgb]{0,0,0}Y}U(s)=L_{T}(s)\grave{Y}_{\color[rgb]{0,0,0}N}(s)$ (18)
Note that $\grave{P}_{m}(s)$ is a minimal polynomial basis and the right-hand side of (18) is pure polynomial, so based on Adapted Forney’s Theorem, $\Delta_{\color[rgb]{0,0,0}Y}U(s)$ must be polynomial. It means that $\Delta_{\color[rgb]{0,0,0}Y}$ captures all poles of $U(s)$. Moreover, non-singular transformations of $\grave{P}(s)$ preserves the invariant zeros [28] so that $y^{*}_{z}\grave{P}_{m}(z)=0u^{*}_{z}=0$. Left multiplying $y^{*}_{z}$ to (18) yields
$\underbrace{y^{*}_{z}\grave{P}_{m}(z)}_{=0}\Delta_{\color[rgb]{0,0,0}Y}(z)U(z)=\underbrace{y^{*}_{z}L_{T}(z)}_{\neq 0}\grave{Y}_{\color[rgb]{0,0,0}N}(z)$ (19)
Since the left-hand side of (19) becomes zero, $\det(L_{T}{\color[rgb]{0,0,0}}){\color[rgb]{0,0,0}=\text{constant}}\neq 0$ (by polynomial-unimodularity), and $1/\Delta_{\color[rgb]{0,0,0}Y}$ is stable because $U(s)\in\mathcal{RH}_{\infty}$ is assumed in this direction, the only way to satisfy the equality is $\grave{Y}_{\color[rgb]{0,0,0}N}(z)=0$ implying $Y(z)=0$.
@@
$z^{+}(s)=\prod_{z\in Z_{R}\bigcap\mathbb{R}}(s-z)^{m_{z}}\prod_{z\in Z_{R}\bigcap(\mathbb{C}\setminus\mathbb{R})}(s-z)^{m_{z}}(s-{z}^{*})^{m_{{z}^{*}}}$
@@
While the algebraic equation is inspired by Exact Model Matching, we explicitly define necessary and sufficient conditions on the system’s output, applicable even to non-minimum phase and singular systems. Specifically, the condition $Y(s)\in\operatorname*{Im}_{\mathbb{R}(s)}(P)$ defines the set of reachable outputs, while $Y(z)=0$ for all $z\in Z_{R}$ defines the set of outputs for which a corresponding stable input exists. Relaxing either condition results in losing the property ‘$=0$” over the $e(t)$.

### v4

@@ Abstract
In this paper, we establish necessary and sufficient conditions for stable inversion, addressing challenges in non-minimum phase, non-square, and singular systems. An $\mathcal{H}_{\infty}$-based algebraic approximation is introduced for near-perfect tracking without preview. Additionally, we propose a novel robust control strategy combining the nominal model with dual feedforward control to form a feedback structure. Numerical comparison demonstrates the approach's effectiveness.
@@ I Introduction
The study of stable inversion began with Brockett's introduction of functional reproducibility in 1965 [2]. Silverman extended these concepts to multivariable systems in 1969 [3], followed by geometric formulations introduced by Basile and Marro [4] and applications in reduced-order control by Moylan [5]. The stable inversion in discrete-time systems was first addressed by Tomizuka [6], while Hunt explored non-causal inversion techniques [7].
@@ II-A Algebraic Preliminaries
where $U_{R}(s)\in\mathbb{R}[s]^{n_{y}\times n_{y}}$ and $V_{R}(s)\in\mathbb{R}[s]^{n_{u}\times n_{u}}$ are $\mathcal{R}$-Unimodular matrices, $\Lambda(s)=\text{diag}[d_{1}(s)\ \dots\ d_{r}(s)]$ with unique monic $d_{i}(s)\in\mathcal{R}$ such that $d_{i}$ divides $d_{i+1}$. The matrix $S_{m}(s)\in\mathbb{R}[s]^{n_{y}\times n_{u}}$ is called the Smith Normal form of $\grave{P}$.
Let $V$ be a vector space over the field $\mathbb{R}(s)$ with dimension $k$, consisting of $n$-tuples such that a basis of vector polynomials can always be found for $V$ [page 22, [22]]. A minimal basis of $V$ is defined as a $k\times n$ polynomial matrix $P_{m}$ [20]. Adapted Forney's Theorem [20, Section 3(4.)] states that if $y=xP_{m}$ is a polynomial $n$-tuple, then $x$ must be a polynomial $k$-tuple.
For the rank notation, akin to [23], the rank of $P$ is defined as the maximum size of any linearly independent subset of its columns in the field $\mathbb{R}(s)$, denoted by $\text{rank}_{\mathbb{R}(s)}(P)$, where $\text{rank}_{\mathbb{R}(s)}(P)\neq\text{rank}_{\mathbb{R}}(P)$.
Suppose $\text{rank}_{\mathbb{R}(s)}(P)=r$, where $1\leq r\leq\min(n_{u},n_{y})$. Define $J=\{j_{1},\dots,j_{r}\}\subseteq\{1,\dots,n_{u}\}$ as an ordered indexed set corresponding to $P(s)$'s linearly independent columns. We then define the matrix $L$ as
@@
Some further notations throughout the paper are: $\|(\cdot)(t)\|$ denotes the Euclidean norm, $\|(\cdot)(t)\|_{\infty}\triangleq\text{ess sup}_{t\geq 0}\|(\cdot)(t)\|$, any complex number can be expressed as $\Re(\cdot)\pm j\Im(\cdot)\in\mathbb{C}$, $\sigma$ represents singular values. Time domain square-integrable functions are denoted by $\mathcal{L}_{2}(-\infty,\infty)$. Its causal subset is given by $\mathcal{L}_{2}[0,\infty)$. For the frequency domain, including at $\infty$, $\mathcal{L}_{2}(j\mathbb{R})$ represents square-integrable functions on $j\mathbb{R}$, and $\mathcal{L}_{\infty}(j\mathbb{R})$ denotes bounded functions on $\Re(s)=0$. All these functions in Lebesgue spaces may be either matrix-valued or scalar. Then, $\mathcal{R}\mathcal{H}_{\infty}$ denotes the set of real rational $\mathcal{L}_{\infty}(j\mathbb{R})$ functions analytic in $\Re(s)>0$, $(\cdot)^{\dagger}$ represents pseudo-inverse yielding identity matrix under multiplication, $(\cdot)^{*}$ denotes complex conjugates transpose, and $\langle.,.\rangle$ denotes the inner product. Among various definitions available for multivariable zeros of a Transfer Function Matrix (TFM), $P(s)$, we adopt the definition given in [24, chapter 4.5.3] as
@@ II-B Problem Statement
where $A\in\mathbb{R}^{n\times n}$, $B\in\mathbb{R}^{n\times n_{u}}$, $C\in\mathbb{R}^{n_{y}\times n}$, $x,x_{0}\in\mathbb{X}\subseteq\mathbb{R}^{n},x(t)\in\mathcal{L}_{2}[0,\infty),y\in\mathbb{Y}\subset\mathbb{R}^{n_{y}},y(t)\in\mathcal{L}_{2}[0,\infty),u\in\mathbb{U}\subset\mathbb{R}^{n_{u}}$, and $u(t)\in\mathcal{L}_{2}[0,\infty)$. Moreover, the TFM representation of $\Xi(u,0)\triangleq C(sI-A)^{-1}B=P(s)$ and since all real rational strictly proper transfer matrices with no poles on the imaginary axis form $\mathcal{RL}_{2}(j\mathbb{R})\subset\mathcal{L}_{2}(j\mathbb{R})$ [25, page 48], $P(s)$ belongs to $\mathcal{L}_{2}(j\mathbb{R})\cap\mathbb{R}(s)^{n_{y}\times n_{u}}$. The system, $\Xi$, is: P1) a non-minimum phase s.t. $Z_{R}\neq\emptyset$, P2) either $n_{y}=n_{u}$ (square) or $n_{y}\neq n_{u}$ (nonsquare), and causal ($t>0$).
@@
Under Assumption 1, let $\Xi$ denote the forward system given in (6), and let $y_{o}(t)$ be an observed output trajectory. We define the right inverse $\Xi_{i}^{\Gamma}:\mathbb{Y}\times\mathbb{X}\to\mathbb{U}$, which, given $y_{o}(t)$ and a prescribed inverse system's initial states, $\bar{x}_{0}\in\mathbb{X}$, produces the control input
$u_{inv}(t)=\Xi_{i}^{\Gamma}(y_{o}(t),\bar{x}_{0}),\quad\|u_{inv}(t)\|_{\infty}<\infty.$ (7)
@@
The terminology here (cf. [26]) also lets $y_{o}(t)$ denote a desired trajectory [10]. With Assumption 1 and $\|e\|_{\infty}<\infty$, (8) is Lyapunov-stable. If exact inversion ($\alpha=\varphi=0$) is unattainable, we can employ a robust stabilizing approximate inverse—one viable choice among the many approximations acknowledged in [11]—with an irreducible error $\varphi$.
@@
As is typical in exact stable inversion, we initially assume that $(A,B,C)$ in (6) and the initial states $x_{0}$ and $\bar{x}_{0}$ are known. Section III-C relaxes this assumption by dropping the requirement on $\bar{x}_{0}$. Section III-D goes further by allowing $x_{0}$ to be unknown and by also incorporating model uncertainty, resulting in an approximate solution. For further details on initial states, see [12].
To solve Problem 1.a algebraically, define $_{ic}(t)\triangleq\Xi(0,x_{0}), yy$$\triangleq y_{o}-y_{ic}$, $u_{inv}^{(0)}(t)\triangleq\Xi_{i}^{\Gamma}(0,\bar{x}_{0})$, and $u(t)\triangleq\Xi_{i}^{\Gamma}(y(t),0)$ to represent the trajectories and control inputs for known or zero initial states. Then, Remark 2 lets us redefine (7) as $u_{inv}(t)\triangleq\Xi_{i}^{\Gamma}(y_{o}(t),\bar{x}_{0})-u_{inv}^{(0)}(t)$. Applying the Laplace transform (see Paley–Wiener Th. [27, p. 104] for existence) yields
$\mathscr{L}\{y_{o}(t)-{y}_{ic}(t)\}={C}(s{I}-{A})^{-1}{B}U(s)=Y(s)$ (10)
@@
From this point on, similar to [19, Theorem 8.5.2], Problem 1.a reduces to algebraically solving the rational matrix equation where $\|u(t)\|_{\infty}<\infty\iff\|u_{inv}(t)\|_{\infty}<\infty$.
@@ III-A Right Inverse
where $U(s)\in\mathbb{R}(s)^{n_{u}}$ and $\grave{Y}(s)\in\mathbb{R}(s)^{n_{y}}$. Then, using the invariant factor theorem given by (II-A) on $\grave{P}(s)$ in (13), we have
$U(s)$ $=[\hat{V}_{R1}\ \hat{V}_{R2}]\begin{bmatrix}\Lambda^{-1}(s)&\textbf{0}_{r\times(n_{y}-r)}\\ \textbf{0}_{(n_{u}-r)\times r}&\textbf{0}_{(n_{u}-r)\times(n_{y}-r)}\end{bmatrix}\begin{bmatrix}\hat{U}_{R1}\\ \hat{U}_{R2}\end{bmatrix}\grave{Y}(s)$
@@
$=\Biggr{(}[\hat{V}_{R1}\ \hat{V}_{R2}]\begin{bmatrix}\Lambda^{-1}(s)&\textbf{0}_{r\times(n_{y}-r)}\\ \textbf{0}_{(n_{u}-r)\times r}&\textbf{0}_{(n_{u}-r)\times(n_{y}-r)}\end{bmatrix}\begin{bmatrix}\hat{U}_{R1}\\ \hat{U}_{R2}\end{bmatrix}$
@@ III-B Algebraic Necessary and Sufficient Conditions
(2:) Note that $(I_{n_{u}}-\hat{V}_{R1}{V}_{R1})$ in (III-A) is pure polynomial and $\kappa\in\mathcal{R}\mathcal{H}_{\infty}$. Therefore, $(I_{n_{u}}-\hat{V}_{R1}{V}_{R1})\kappa$ does not contribute any unstable solution(s) to $U(s)$. So, considering only the $\hat{V}_{R1}(s)\Lambda^{-1}(s)\hat{U}_{R1}(s)\grave{Y}(s)$ is enough for stability. Let's consider two cases on $\grave{P}(s)$; C1: $\operatorname*{rank}_{\mathbb{R}(s)}(\grave{P})=\min(n_{u},n_{y})$, C2: $\operatorname*{rank}_{\mathbb{R}(s)}(\grave{P})<\min(n_{u},n_{y})$. For C2, re-writing (III-A) and ignoring $(I_{n_{u}}-\hat{V}_{R1}{V}_{R1})\kappa$, we get
@@
where there always exist $\hat{V}^{\ell}_{R1}$ satisfying $\hat{V}^{\ell}_{R1}\hat{V}_{R1}=I_{r}$. Now assume that either $\grave{P}(s)$ or $U_{R1}\Lambda\hat{V}_{R1}^{\ell}$ is not a minimal basis. The minimal basis $\grave{P}_{m}(s)$ of an $n_{y}$-dimensional vector space of $n_{u}$-tuples over $\mathbb{R}[s]$ for (13) is given by $\grave{P}_{m}(s)=L_{T}(s)\grave{P}(s)$ for condition C1, and $\grave{P}_{m}(s)=L_{T}(s)U_{R1}\Lambda\hat{V}_{R1}^{\ell}$ for condition C2. Here, the $\mathcal{R}$-Unimodular transformation $L_{T}(s)\in\mathbb{R}[s]^{n_{y}\times n_{y}}$, which yields the minimal basis, exists with constant (degree zero) determinant (see [20, Remark 2]). Then, left multiplying $L_{T}(s)$ to (13) or (16) results in
@@
Now define $\grave{Y}(s)\triangleq\frac{1}{\Delta_{Y}}{\grave{Y}_{N}}(s)$ such that $\grave{Y}_{N}(s)\in\mathbb{R}[s]^{n_{y}}$, and the scalar $\Delta_{Y}\in\mathbb{R}[s]$. So, multiplying (17) by $\Delta_{Y}$ yields
$\grave{P}_{m}(s)\Delta_{Y}U(s)=L_{T}(s)\grave{Y}_{N}(s)$ (18)
Note that $\grave{P}_{m}(s)$ is a minimal polynomial basis and the right-hand side of (18) is pure polynomial, so based on Adapted Forney's Theorem, $\Delta_{Y}U(s)$ must be polynomial. It means that $\Delta_{Y}$ captures all poles of $U(s)$. Moreover, non-singular transformations of $\grave{P}(s)$ preserves the invariant zeros [28] so that $y^{*}_{z}\grave{P}_{m}(z)=0u^{*}_{z}=0$. Left multiplying $y^{*}_{z}$ to (18) yields
$\underbrace{y^{*}_{z}\grave{P}_{m}(z)}_{=0}\Delta_{Y}(z)U(z)=\underbrace{y^{*}_{z}L_{T}(z)}_{\neq 0}\grave{Y}_{N}(z)$ (19)
Since the left-hand side of (19) becomes zero, $\det(L_{T})=\text{constant}\neq 0$ (by polynomial-unimodularity), and $1/\Delta_{Y}$ is stable because $U(s)\in\mathcal{RH}_{\infty}$ is assumed in this direction, the only way to satisfy the equality is $\grave{Y}_{N}(z)=0$ implying $Y(z)=0$.
@@
$z^{+}(s)=\prod_{z\in Z_{R}\bigcap\mathbb{R}}(s-z)^{m_{z}}\prod_{z\in Z_{R}\bigcap(\mathbb{C}\setminus\mathbb{R})}(s-z)^{m_{z}^{\ast}}(s-{z}^{\ast})^{m_{z}^{\ast}}$
@@
While the algebraic equation is inspired by Exact Model Matching, we explicitly define necessary and sufficient conditions on the system's output, applicable even to non-minimum phase and singular systems. Specifically, the condition $Y(s)\in\operatorname*{Im}_{\mathbb{R}(s)}(P)$ defines the set of reachable outputs, while $Y(z)=0$ for all $z\in Z_{R}$ defines the set of outputs for which a corresponding stable input exists. Relaxing either condition results in losing the property `=0=0'' over the $e(t)$.